# An Efficient Row-Based Sparse Fine-Tuning

### v1

@@ C.5 Cache Benefit
$r=32$, random 43.34 76.43 77.83 57.18 31.4 78.13 62.45 45.80 69.30 60.21
$r=32$, $\ell^{2}$ 43.34 77.43 77.80 57.06 32.2 78.02 63.18 46.52 69.14 60.52
$r=32$, ZOTaylor 43.43 77.44 77.49 57.79 32.0 78.35 64.26 46.67 68.90 60.70
$r=64$, random 43.17 76.39 77.86 57.20 31.2 78.07 62.45 45.91 69.06 60.15
$r=64$, $\ell^{2}$ 43.69 77.30 77.83 57.19 32.2 78.07 63.54 47.03 69.22 60.67
$r=64$, ZOTaylor 44.28 76.89 77.49 57.85 32.0 78.67 65.34 46.72 68.98 60.92
$r=128$, random 43.34 76.30 77.80 57.22 31.2 78.07 63.18 45.96 69.22 60.25
$r=128$, $\ell^{2}$ 43.60 77.26 77.77 57.47 32.6 78.07 64.98 46.67 69.30 60.86
$r=128$, ZOTaylor 44.11 77.15 77.43 57.71 32.4 78.35 65.70 46.67 68.90 60.94
$r=32$, FA, random 43.26 76.39 77.28 57.14 31.4 78.07 62.82 46.06 69.30 60.19
$r=32$, FA, $\ell^{2}$ 43.69 77.30 77.55 57.14 31.6 78.02 64.62 46.88 69.38 60.69
$r=32$, FA, ZOTaylor 44.20 77.61 77.22 57.41 31.8 78.07 65.70 46.72 68.90 60.85
$r=64$, FA, random 43.34 76.30 77.52 57.18 31.4 78.07 62.82 46.01 69.14 60.20
$r=64$, FA, $\ell^{2}$ 43.77 77.26 77.65 57.17 31.8 78.07 65.70 46.78 69.22 60.82
$r=64$, FA, ZOTaylor 43.94 77.44 77.71 57.43 32.6 77.91 66.06 46.42 68.75 60.92
$r=128$, FA, random 43.34 76.43 77.46 57.19 31.4 78.02 64.26 46.01 69.06 60.35
$r=128$, FA, $\ell^{2}$ 43.94 77.22 77.83 57.11 32.0 78.18 65.70 46.47 69.38 60.87
$r=128$, FA, ZOTaylor 44.45 77.82 77.74 57.41 31.4 77.80 66.06 46.57 68.59 60.87
@@
$r=32$, random 50.51 80.26 81.59 60.20 34.6 79.54 68.95 47.08 72.53 63.92
$r=32$, $\ell^{2}$ 51.19 80.81 81.10 60.21 34.4 79.60 70.40 47.19 73.01 64.21
$r=32$, ZOTaylor 50.94 81.61 82.42 60.17 35.6 79.60 70.40 47.19 73.40 64.59
$r=64$, random 50.51 80.09 81.44 60.22 34.6 79.60 69.31 47.24 72.69 63.97
$r=64$, $\ell^{2}$ 51.96 81.27 81.31 60.18 35.2 79.76 70.04 47.80 73.09 64.51
$r=64$, ZOTaylor 52.13 81.14 82.32 60.19 35.6 79.82 70.04 47.03 73.01 64.59
$r=128$, random 50.43 80.13 81.47 60.19 34.6 79.60 69.68 47.24 72.69 64.00
$r=128$, $\ell^{2}$ 52.47 81.10 81.28 60.29 34.6 79.76 70.04 47.75 73.24 64.50
$r=128$, ZOTaylor 52.22 81.99 82.54 60.13 34.6 79.76 70.04 47.59 73.16 64.67
$r=32$, FA, random 50.43 80.05 81.47 60.19 34.6 79.65 69.68 47.03 72.85 63.99
$r=32$, FA, $\ell^{2}$ 50.60 80.85 81.19 60.22 34.2 79.54 70.76 47.24 73.01 64.18
$r=32$, FA, ZOTaylor 51.02 81.31 82.05 60.21 35.2 79.71 70.40 47.03 73.24 64.46
$r=64$, FA, random 50.26 80.13 81.53 60.20 34.4 79.54 68.95 47.03 72.93 63.89
$r=64$, FA, $\ell^{2}$ 51.79 81.06 81.22 60.20 35.2 79.65 71.48 47.24 73.16 64.56
$r=64$, FA, ZOTaylor 51.71 81.31 82.05 60.22 34.8 79.65 69.68 47.29 72.77 64.39
$r=128$, FA, random 50.26 80.09 81.47 60.22 34.6 79.71 69.31 47.08 72.85 63.95
$r=128$, FA, $\ell^{2}$ 52.22 81.19 81.35 60.20 34.2 79.71 69.31 47.13 73.01 64.26
$r=128$, FA, ZOTaylor 52.22 81.23 82.17 60.19 34.4 79.43 69.68 47.13 73.01 64.38
Table 13: Similar to Table 6, this is the full table with different rank settings. We also present the results of freezing-attention blocks in this table.
C.6 Rank Settings
@@
Table 13 is the full table of importance evaluation for Llama2 and Llama3.
C.7 Benefit of Freezing Attention Blocks
Model, ft setting mem #param ARC-c ARC-e BoolQ HS OBQA PIQA rte SIQA WG Avg
Llama2(7B), LoRA, $r=16$ 21.64GB 40.0M(0.59%) 44.71 76.89 77.49 57.94 32.2 78.73 60.65 47.59 68.75 60.55
Llama2(7B), FA-LoRA, $r=32$ 16.56GB 46.4M(0.69%) 44.03 77.48 77.61 57.40 30.4 77.80 65.70 46.98 68.98 60.71
Llama2(7B), LoRA, $r=32$ 22.21GB 80.0M(1.19%) 44.28 76.89 77.37 57.61 32.0 78.45 64.62 47.90 69.14 60.92
Llama2(7B), FA-LoRA, $r=64$ 17.25GB 92.8M(1.38%) 43.77 77.57 77.74 57.45 31.0 77.86 66.06 47.13 69.06 60.85
Llama2(7B), SPruFT, $r=32$ 15.57GB 36.4M(0.54%) 43.34 77.43 77.80 57.06 32.2 78.02 63.18 46.52 69.14 60.52
Llama2(7B), FA-SPruFT, $r=64$ 14.48GB 39.3M(0.58%) 43.77 77.26 77.65 57.17 31.8 78.07 65.70 46.78 69.22 60.82
Llama2(7B), SPruFT, $r=64$ 16.20GB 72.9M(1.08%) 43.69 77.30 77.83 57.19 32.2 78.07 63.54 47.03 69.22 60.67
Llama2(7B), FA-SPruFT, $r=128$ 15.21GB 78.6M(1.17%) 43.94 77.22 77.83 57.11 32.0 78.18 65.70 46.47 69.38 60.87
Llama3(8B), LoRA, $r=16$ 28.86GB 41.9M(0.52%) 53.50 81.44 82.35 60.61 34.2 79.87 69.31 48.00 73.56 64.76
Llama3(8B), FA-LoRA, $r=32$ 23.89GB 56.6M(0.71%) 52.22 81.61 82.35 60.26 35.0 79.98 69.68 48.41 73.80 64.81
Llama3(8B), LoRA, $r=32$ 29.37GB 83.9M(1.04%) 53.33 81.86 82.20 60.65 34.0 79.87 68.23 49.03 73.72 64.76
Llama3(8B), FA-LoRA, $r=64$ 24.55GB 113.2M(1.41%) 52.47 81.36 82.23 60.17 35.0 79.76 70.04 48.31 73.56 64.77
Llama3(8B), SPruFT, $r=32$ 22.62GB 39.8M(0.50%) 51.19 80.81 81.10 60.21 34.4 79.60 70.40 47.19 73.01 64.21
Llama3(8B), FA-SPruFT, $r=64$ 21.62GB 46.1M(0.57%) 51.79 81.06 81.22 60.20 35.2 79.65 71.48 47.24 73.16 64.56
Llama3(8B), SPruFT, $r=64$ 23.23GB 79.7M(0.99%) 51.96 81.27 81.31 60.18 35.2 79.76 70.04 47.80 73.09 64.51
Llama3(8B), FA-SPruFT, $r=128$ 22.41GB 92.3M(1.15%) 52.22 81.19 81.35 60.20 34.2 79.71 69.31 47.13 73.01 64.26
Table 14: Same results of Table 12 with a reordering of the rows. This table is for comparing fine-tuning all linear layers with fine-tuning only the MLP layers.
@@ C.7 Benefit of Freezing Attention Blocks
Appendix D Details of Datasets
D.1 Vision Benchmarks
@@
D.2 General Language Understanding Evaluation Benchmark (GLUE)
@@
RTE: The Recognizing Textual Entailment (RTE) datasets are a series of challenges that evaluate models’ ability to determine whether a premise can entail a given hypothesis (Dagan et al., 2006; Bar-Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009). The data are constructed based on the texts from Wikipedia and news. The datasets have been used to evaluate the performance of both traditional language models and the state-of-the-art LLMs.
@@
D.3 Text-Generation Datasets
Stanford Alpaca: Alpaca is an instruction dataset designed for instruction training of pre-trained language models (Taori et al., 2023). It contains 52002 instruction-response pairs generated by OpenAI’s text-davinci-003 engine or written by humans. Note that there is only a training split in this dataset. Models fine-tuned on Alpaca are often evaluated by other tasks like “EleutherAI LM Harness”.

### v2

@@ D.5 Cache Benefit
$r=32$, random 72.2 70.8 65.6 77.0 55.4 71.0 77.0 69.2 69.80
$r=32$, $\ell^{2}$ 78.2 73.6 67.4 79.4 58.8 69.0 82.2 69.4 72.25
$r=32$, ZOTaylor 74.4 72.0 66.6 80.4 63.4 71.8 76.6 73.6 72.35
$r=64$, random 71.0 68.0 62.8 74.2 63.6 66.6 78.0 71.4 69.45
$r=64$, $\ell^{2}$ 78.0 73.8 66.4 83.4 59.6 69.4 79.0 71.6 72.65
$r=64$, ZOTaylor 78.2 71.0 68.4 80.8 62.0 69.6 81.0 71.4 72.80
$r=128$, random 80.0 70.6 69.8 68.2 57.4 67.6 74.6 70.2 69.80
$r=128$, $\ell^{2}$ 80.0 75.2 67.6 85.0 63.4 70.8 82.4 71.8 74.53
$r=128$, ZOTaylor 79.4 74.4 67.0 82.4 65.0 72.4 83.0 71.2 74.35
$r=32^{{\ddagger}}$, random 78.0 71.2 64.8 72.6 55.8 67.8 76.4 69.8 69.55
$r=32^{{\ddagger}}$, $\ell^{2}$ 82.6 73.0 64.4 80.0 61.4 64.8 80.4 70.2 72.10
$r=32^{{\ddagger}}$, ZOTaylor 77.4 71.2 65.6 78.0 59.4 68.8 79.8 69.0 71.15
$r=64^{{\ddagger}}$, random 75.8 71.0 64.4 70.0 58.8 71.2 79.8 72.0 70.38
$r=64^{{\ddagger}}$, $\ell^{2}$ 78.2 75.8 66.4 81.4 59.2 67.8 78.8 72.0 72.45
$r=64^{{\ddagger}}$, ZOTaylor 81.4 71.4 67.0 82.2 61.6 71.2 80.0 72.0 73.34
$r=128^{{\ddagger}}$, random 75.8 75.2 67.2 79.2 64.0 69.6 77.0 68.6 72.08
$r=128^{{\ddagger}}$, $\ell^{2}$ 81.4 75.0 67.2 80.4 65.0 69.2 79.8 73.8 73.98
$r=128^{{\ddagger}}$, ZOTaylor 81.0 74.4 69.0 81.0 64.6 70.6 81.2 72.8 74.30
@@
$r=32$, random 83.8 78.4 59.2 86.4 69.8 79.4 88.4 76.6 77.75
$r=32$, $\ell^{2}$ 81.2 82.2 68.2 85.4 63.4 79.4 87.6 80.8 78.53
$r=32$, ZOTaylor 81.6 78.4 70.4 85.0 63.2 79.2 88.2 84.6 78.83
$r=64$, random 82.6 73.4 68.8 72.6 63.4 74.0 81.6 75.0 73.93
$r=64$, $\ell^{2}$ 83.2 80.6 69.4 86.0 60.6 78.4 83.4 81.4 77.88
$r=64$, ZOTaylor 81.8 78.0 68.4 85.4 64.6 77.6 85.8 82.4 78.00
$r=128$, random 84.2 77.4 70.2 72.0 72.4 72.8 84.0 75.6 76.08
$r=128$, $\ell^{2}$ 87.6 77.4 71.4 85.4 70.2 79.8 90.8 81.8 80.55
$r=128$, ZOTaylor 89.0 78.8 70.6 86.2 69.4 80.4 92.0 83.8 81.28
$r=32^{{\ddagger}}$, random 81.6 80.0 69.2 87.6 69.8 82.4 89.6 85.0 80.65
$r=32^{{\ddagger}}$, $\ell^{2}$ 83.2 82.6 66.6 86.0 65.6 82.4 89.8 80.4 79.58
$r=32^{{\ddagger}}$, ZOTaylor 86.2 77.6 72.2 84.0 65.6 78.8 90.0 86.4 80.10
$r=64^{{\ddagger}}$, random 81.2 80.2 61.8 82.8 70.6 79.6 89.2 82.0 78.43
$r=64^{{\ddagger}}$, $\ell^{2}$ 84.4 78.0 71.0 83.2 69.0 79.0 84.4 83.4 79.05
$r=64^{{\ddagger}}$, ZOTaylor 84.6 79.0 72.4 88.0 61.2 78.4 88.8 86.8 79.90
$r=128^{{\ddagger}}$, random 82.8 75.2 68.8 83.8 67.8 78.2 83.0 78.0 77.20
$r=128^{{\ddagger}}$, $\ell^{2}$ 89.4 79.0 71.2 86.2 71.8 83.0 86.8 76.2 80.45
$r=128^{{\ddagger}}$, ZOTaylor 90.6 77.0 71.0 84.0 72.6 80.0 92.6 86.6 81.80
Table 16: Importance evaluation for Llama2 and Llama3 on MT-Bench. We also present the results of freezing queue-, key-, and value-projection in this table (). Bold indicates the best result on the same task.
Model, ft setting Coding Extraction Humanities Math Reasoning Roleplay Stem Writing Avg
Llama2(7B), SPruFT
$r=128$, random 0.67 3.44 5.11 1.67 3.50 4.95 4.41 3.89 3.45
$r=128$, $\ell$ 1.82 2.55 4.80 2.08 4.07 4.79 4.50 3.53 3.52
$r=128$, ZOTaylor 1.42 2.63 5.16 2.36 3.67 5.20 4.39 4.11 3.62
$r=128^{{\ddagger}}$, random 2.31 2.90 4.75 2.77 2.94 4.61 4.35 4.15 3.60
$r=128^{{\ddagger}}$, $\ell$ 1.08 2.10 4.60 1.15 3.80 4.22 4.33 3.42 3.09
$r=128^{{\ddagger}}$, ZOTaylor 2.44 2.63 5.20 2.21 2.93 5.00 4.94 3.95 3.66
Llama3(8B), SPruFT
$r=128$, random 3.22 3.74 4.72 3.33 3.50 5.20 6.38 5.11 4.40
$r=128$, $\ell$ 3.88 5.11 6.11 3.83 4.21 5.35 6.44 5.40 5.04
$r=128$, ZOTaylor 4.13 4.78 6.89 6.33 4.08 5.95 5.39 5.16 5.34
$r=128^{{\ddagger}}$, random 3.56 5.06 5.68 4.69 4.00 5.26 6.17 6.00 5.05
$r=128^{{\ddagger}}$, $\ell$ 4.75 4.78 5.16 3.67 3.31 6.25 6.06 5.00 4.87
$r=128^{{\ddagger}}$, ZOTaylor 4.13 5.38 6.05 4.79 5.00 5.22 5.88 5.21 5.21
D.6 Rank Settings
@@
Table 15 and Table 16 are the full tables of importance evaluation for Llama2 and Llama3.
D.7 Benefit of Freezing Attention Blocks
Table 17: Same results of Table 14 with a reordering of the rows. This table is for comparing fine-tuning all linear layers with freezing queue-, key-, and value-projection.
Model, ft setting mem #param BoolQ PIQA SIQA HS WG ARC-c ARC-e OBQA Avg
Llama2(7B), LoRA, $r=16$ 21.64GB 40.0M(0.59%) 76.4 71.4 69.4 81.2 59.8 67.0 81.4 73.2 72.48
Llama2(7B), LoRA, $r=32$ 17.95GB 54.8M(0.81%) 75.6 76.0 68.6 79.2 63.6 67.2 82.0 71.0 72.90
Llama2(7B), LoRA, $r=32$ 22.21GB 80.0M(1.19%) 74.8 74.2 71.4 78.4 58.6 67.0 82.2 69.6 72.03
Llama2(7B), LoRA, $r=64$ 18.81GB 109.6M(1.63%) 79.6 73.8 67.0 80.4 62.6 69.2 81.0 70.4 73.00
Llama2(7B), SPruFT, $r=32$ 15.57GB 36.4M(0.54%) 78.2 73.6 67.4 79.4 58.8 69.0 82.2 69.4 72.25
Llama2(7B), SPruFT, $r=64$ 14.67GB 47.7M(0.71%) 78.2 75.8 66.4 81.4 59.2 67.8 78.8 72.0 72.45
Llama2(7B), SPruFT, $r=64$ 16.20GB 72.9M(1.08%) 78.0 73.8 66.4 83.4 59.6 69.4 79.0 71.6 72.65
Llama2(7B), SPruFT, $r=128$ 15.58GB 95.4M(1.42%) 81.4 75.0 67.2 80.4 65.0 69.2 79.8 73.8 73.98
Llama3(8B), LoRA, $r=16$ 28.86GB 41.9M(0.52%) 85.2 80.8 68.4 81.8 69.0 79.4 90.0 77.0 78.95
Llama3(8B), LoRA, $r=32$ 25.28GB 65.0M(0.81%) 83.0 68.8 71.4 85.2 66.2 81.0 91.6 79.6 78.35
Llama3(8B), LoRA, $r=32$ 29.37GB 83.9M(1.04%) 85.2 81.8 68.2 87.8 67.0 76.4 89.2 80.4 79.50
Llama3(8B), LoRA, $r=64$ 26.04GB 130.0M(1.62%) 86.2 81.6 67.8 81.8 66.0 73.8 86.2 78.2 77.73
Llama3(8B), SPruFT, $r=32$ 22.62GB 39.8M(0.50%) 81.2 82.2 68.2 85.4 63.4 79.4 87.6 80.8 78.53
Llama3(8B), SPruFT, $r=64$ 21.81GB 54.5M(0.68%) 84.4 78.0 71.0 83.2 69.0 79.0 84.4 83.4 79.05
Llama3(8B), SPruFT, $r=64$ 23.23GB 79.7M(0.99%) 83.2 80.6 69.4 86.0 60.6 78.4 83.4 81.4 77.88
Llama3(8B), SPruFT, $r=128$ 22.71GB 109.1M(1.36%) 89.4 79.0 71.2 86.2 71.8 83.0 86.8 76.2 80.45
@@ D.7 Benefit of Freezing Attention Blocks
Appendix E Details of Datasets
E.1 Vision Benchmarks
@@
E.2 General Language Understanding Evaluation Benchmark (GLUE)
@@
RTE: The Recognizing Textual Entailment (RTE) datasets are a series of challenges that evaluate models’ ability to determine whether a premise can entail a given hypothesis (Dagan et al., 2006; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009). The data are constructed based on the texts from Wikipedia and news. The datasets have been used to evaluate the performance of both traditional language models and the state-of-the-art LLMs.
@@
E.3 Text-Generation Datasets
[…]
Stanford Alpaca: Alpaca is an instruction dataset designed for instruction training of pre-trained language models (Taori et al., 2023). It contains 52002 instruction-response pairs generated by OpenAI’s text-davinci-003 engine or written by humans. Note that there is only a training split in this dataset. Models fine-tuned on Alpaca are often evaluated by other tasks like “EleutherAI LM Harness”. Alpaca-GPT is an updated version with the answers generated by GPT-4 (Achiam et al., 2023).